OpenStudy (anonymous):

Will Medal and fan, please help! Use a net to find the surface area of the right triangular prism shown below

OpenStudy (anonymous):

https://lvs.brainhoney.com/Resource/45133303,0/Assets/84734_5508510b/05_07_18.jpg 104 square feet 412 square feet 468 square feet 504 square feet

OpenStudy (anonymous):

@phi @pooja195 @satellite73 @texaschic101 @Directrix @freckles @mathmale @YanaSidlinskiy

OpenStudy (anonymous):

add up the three rectangles area, and the two triangles area

OpenStudy (anonymous):

add up the surface areas of each side now. 3 rectangles 2 triangles

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

rectangles = 8*13 + 8 *12 + 8*5 triangles = 5*12

OpenStudy (anonymous):

300

OpenStudy (anonymous):

yes, 300

OpenStudy (anonymous):

104 square feet 412 square feet 468 square feet 504 square feet